# Hai-Long Sun1,2,3   Da-Wei Zhou1,2  Yang Li3  Shiyin Lu3  Chao Yi1,2 Qing-Guo Chen3  Zhao Xu3  Weihua Luo3  Kaifu Zhang3  De-Chuan Zhan1,2  Han-Jia Ye1,2

### v2

@@ 5.3 Ablation Study
Figure 5: Left: The ablation study of multilingual data and the MoE module using the MMBench benchmark. Middle: The performance of Parrot on a broad range of multimodal tasks compared with existing models. Models with 7B parameters are used for the two experiments. Right: Expert distributions of MoE. We summarize the activated experts during the feed-forward process using Chinese Prompts.
Dataset English Chinese Portuguese Arabic Turkish Russian
LLaVA-1.5-finetune 69.4 66.6 60.3 55.3 52.1 60.7
+ zh 69.2 -0.2   68.6 +2.0   64.1 +3.8   59.1 +3.8   50.9 -1.2   61.6 +0.9
+ zh pt 71.1 +1.7   70.4 +3.8   65.4 +5.1   57.9 +2.6   52.1 +0.0 62.9 +2.2
+ zh pt ar 71.0 +1.6   68.6 +2.0   65.7 +5.4   58.6 +3.3   52.2 +0.1   62.2 +1.5
+ zh pt ar tr 70.4 +1.0   68.7 +2.1   64.9 +4.6   61.2 +5.9   59.7 +7.6   62.0 +1.3
+ zh pt ar tr ru 70.7 +1.3   70.4 +3.8   65.1 +4.8   57.8 +2.5   58.4 +6.3   64.0 +3.3
Ablation study on monolingual fine-tuning datasets. The ablation study presented in Table LABEL:tab:ablation-mono evaluates the performance of different monolingual datasets added incrementally to the baseline dataset LLaVA-1.5-finetune. It highlights the significant impact of adding different multilingual datasets to a baseline model. Each dataset incrementally improves performance in its respective language and, when combined, leads to overall enhanced performance across all evaluated languages. This indicates the robustness and effectiveness of the proposed method in handling multilingual data, making it a scalable solution for multilingual tasks.
@@ 5.4 Visualization of Multilingual Conversations
To enhance the intuitive understanding of the Parrot’s multilingual capability, we prepare a comprehensive case study accompanied by illustrative visuals. For instance, as depicted in Figure 6, our framework demonstrates remarkable multilingual capabilities. This underscores the Parrot’s versatility in navigating different languages and presents its potential in bridging linguistic gaps across diverse domains. Through careful analysis and visualization, we aim to provide a deeper insight into the mechanism driving this capability, illustrating its practical implications and potential applications in real-world scenarios. This visualization serves as a strong indicator of the Parrot’s solid architecture and its exceptional ability to understand, process, and generate multiple languages with remarkable efficiency. More multilingual conversation cases are shown in Appendix E.
@@ 6 Conclusion
This paper addresses the critical challenge of enhancing the multilingual capabilities of MLLMs. We introduce Parrot, a novel method leveraging textual guidance to drive visual token alignment at the language level, thus enabling the transition of English-biased visual embeddings into language-specific ones using an MoE module. Extensive experiments conducted on a newly introduced Massive Multilingual Multimodal Benchmark (MMMB) across six languages demonstrate the state-of-the-art performance of Parrot compared to existing methods, particularly presenting remarkable improvements in Turkish and Arabic. Furthermore, our model exhibits competitive results across a wide range of diverse multimodal benchmarks, emphasizing its efficacy in addressing both multilingual and multimodal challenges. Parrot not only advances the frontier of MLLMs but also underscores the importance of equitable access to technological benefits across linguistic and cultural diversities.
Appendix A Training Datasets in LLaVA
In this section, we analyze the multilingual data in LLaVA [38]. From Table 4(b) and Figure 7, it is evident that during the pre-train stage, LLaVA solely utilizes multimodal image-text pairs data for training, comprising 558K of English data. During the SFT stage, both multimodal and text-only data are incorporated into the training process. Multilingual data appear only in the text-only dataset. Apart from English, the most prominent non-English data is Chinese, amounting to just 3.1K, constituting 0.25% of the total dataset. Therefore, it is evident that LLaVA’s datasets are English-centric and imbalanced. The specific language and abbreviation are as follows: English (en), Chinese (zh), Korean (ko), Spanish (es), French (fr), Japanese (ja), German (de), Portuguese (pt), Traditional Chinese (zh-tw), Italian (it).
@@ Appendix A Training Datasets in LLaVA
(a) The detailed information in two stages.
@@
Appendix B Training Details

### v3

@@ 4.1 Experimental Setup
Figure 7: Left: The ablation study of multilingual data and the MoE module using the MMBench benchmark. Middle: The performance of Parrot on a broad range of multimodal tasks compared with existing models. Right: Expert distributions of MoE. We summarize the activated experts during the feed-forward process using Chinese Prompts.
4.2 Main Results
As shown in Table 1, Parrot-Qwen2-7B achieves state-of-the-art (SOTA) performance across four languages in both the MMBench and MMMB benchmark, with English and Chinese in second place. Figure 7 demonstrates that Parrot excels not only in multilingual capabilities but also in handling complex multimodal tasks, such as MME (Fu et al., 2023), MMStar (Chen et al., 2024), and SEED-Bench (Li et al., 2024a). In Figure 7, we infer Parrot using a Chinese prompt and visualize the expert distributions within the MoE, revealing the dynamic activation of different experts for different languages. Additionally, as illustrated in Figure 4, Parrot achieves competitive performance in existing multilingual benchmarks, utilizing less than 1% of the data compared to other multilingual MLLMs.
To validate the effectiveness of the Parrot architecture, we use a Chinese prompt (translated to English as “please describe the image in detail”), sample 50 images of ‘tiger cat’ from ImageNet, and perform t-SNE (Van der Maaten & Hinton, 2008) visualizations of both visual and textual features in LLaVA and Parrot. As shown in Figure 6, for LLaVA, the Chinese-CLIP-based visual features are significantly closer in high-dimensional space compared to OpenAI-CLIP. However, thanks to the architecture we designed, Parrot bridges the gap with textual features effectively, regardless of the vision encoder used.
4.3 Further Analysis
In this section, we explore several critical questions to comprehensively analyze Parrot.
  • Are all components of Parrot equally effective? As shown in Figure 7, incorporating multilingual data enhances performance across all languages. Additionally, the MoE module contributes significantly to performance improvements, validating the effectiveness of our proposed method.
  • How does Parrot handle the quality and imbalance challenges? Large-scale translated multilingual data, while seemingly abundant, often suffers from quality issues (e.g., translation errors, cultural mismatches, noisy artifacts), particularly for low-resource languages. Even with massive data volumes, low-resource languages often remain underrepresented as high-resource languages dominate training distributions (e.g., 90%+ of tokens in typical datasets). Forcing higher proportions of low-resource data can risk triggering the “curse of multilingualism,” where models sacrifice high-resource language performance to accommodate low-resource languages, as observed in prior work (Conneau et al., 2019; Chang et al., 2023; Blevins et al., 2024). Our approach strategically prioritizes high-quality alignment signals over raw data quantity. This avoids the pitfalls of noisy translation and the imbalance inherent to brute-force multilingual scaling, ensuring stable performance across all languages without compromising high-resource capabilities.
  • How does Parrot perform compared to a naive, translation-based baseline? To assess this, we implement a baseline using the Google Translation API that translates non-English queries to English, processes them, and translates responses back to the original language. As shown in Table 10, the results reveal a “seesaw effect”: while this naive approach yields improvements in certain languages like Chinese, it simultaneously causes performance degradation in others, particularly Russian and Portuguese. This phenomenon highlights the fundamental limitations of relying solely on translation services for addressing multilingualism in multimodal tasks, underscoring the need for more sophisticated approaches like Parrot.
  • Are the scaling laws effective in Parrot? To explore the effectiveness of scaling laws in multilingual settings, we conduct experiments where the multilingual data (excluding Chinese and English) is gradually expanded until it matches the volume of Chinese data (70K). As shown in Table12, the results indicate that Parrot continues to follow the multilingual scaling law. For example, the performance on Portuguese increased by 3.0 points, and Arabic saw a 5.2-point gain. Additionally, Parrot benefits from model size scaling, as shown in Table 12.
  • Do the main performance gains come from the multilingual dataset? As shown in Table 10 and Figure 7, LLaVA shows limited improvements with the addition of multilingual data. In contrast, Parrot achieves substantial gains, significantly outperforming LLaVA. Therefore, it is ensure that the primary performance gains come from the design of Parrot.
@@ 4.4 Visualization of Multilingual Conversations
To enhance the intuitive understanding of the Parrot’s multilingual capability, we prepare a comprehensive case study accompanied by illustrative visuals. For instance, as depicted in Figure 8, our framework demonstrates remarkable multilingual capabilities. This underscores the Parrot’s versatility in navigating different languages and presents its potential in bridging linguistic gaps across diverse domains. Through careful analysis and visualization, we aim to provide a deeper insight into the mechanism driving this capability, illustrating its practical implications and potential applications in real-world scenarios. This visualization serves as a strong indicator of the Parrot’s solid architecture and its exceptional ability to understand, process, and generate multiple languages with remarkable efficiency. More multilingual conversation cases are shown in Appendix F.
@@ 5 Conclusion
This paper addresses the critical challenge of improving the multilingual capabilities of MLLMs and investigates the misalignment of visual features across languages. We introduce Parrot, a novel approach that leverages textual guidance to align visual tokens at the language level, enabling the conversion of English-biased visual embeddings into language-specific ones through an MoE module. Extensive experiments conducted on a newly introduced benchmark, the Massive Multilingual Multimodal Benchmark (MMMB), across six languages demonstrate that Parrot achieves state-of-the-art performance compared to existing methods, with particularly notable improvements in Turkish and Arabic. Parrot not only advances the frontier of MLLMs but also highlights the importance of equitable access to technological benefits across linguistic and cultural contexts.
Acknowledgments
This work is partially supported by National Key R&D Program of China (2024YFE0202800), NSFC (62376118, 62476123), Key Program of Jiangsu Science Foundation (BK20243012), Fundamental Research Funds for the Central Universities (2024300373, 14380021), the AI & AI for Science Project of Nanjing University, Collaborative Innovation Center of Novel Software Technology and Industrialization.
Impact Statement
Parrot leveraging MoE to enhance multilingual alignment presents a positive social impact by promoting linguistic diversity and inclusivity. To address the challenge of the imbalanced language data in SFT datasets and improve non-English visual tokens alignment, this approach contributes to breaking language barriers and facilitating cross-cultural communication, thereby fostering understanding and collaboration across diverse linguistic communities. Additionally, the creation of the Massive Multilingual Multimodal Benchmark (MMMB) fills a crucial gap in evaluating multilingual capabilities, enabling researchers to assess and improve upon models’ performance across different languages and cultures. This paper presents work whose goal is to advance the field of Machine Learning and the potential societal impact of this work is largely positive.
Appendix A The Details of Training Datasets
In this section, we analyze the multilingual data in LLaVA (Liu et al., 2023b). From Table 2(b) and Figure 10, it is evident that during the pre-train stage, LLaVA solely utilizes multimodal image-text pairs data for training, comprising 558K of English data. During the SFT stage, both multimodal and text-only data are incorporated into the training process. Multilingual data appear only in the text-only dataset. Apart from English, the most prominent non-English data is Chinese, amounting to just 3.1K, constituting 0.25% of the total dataset. Therefore, it is evident that LLaVA’s datasets are English-centric and imbalanced. The specific language and abbreviation are as follows: English (en), Chinese (zh), Korean (ko), Spanish (es), French (fr), Japanese (ja), German (de), Portuguese (pt), Traditional Chinese (zh-tw), Italian (it).
@@ Appendix A The Details of Training Datasets
(a) The language information in two stages.
@@
Figure 10: An example of circular evaluation strategy.
Appendix B Related Work
Multimodal Large Language Models. The domain of MLLMs has witnessed significant advances, particularly in the enhancement of visual and language processing. Current MLLMs are usually a combination of visual encoders (Radford et al., 2021; Sun et al., 2023; Fang et al., 2023; Zhang et al., 2022; Oquab et al., 2023; Zhai et al., 2023), LLMs, and fusion modules. Innovations like Flamingo (Alayrac et al., 2022) have advanced visual representation by integrating a Perceiver Resampler with vision encoders. BLIP-2 (Li et al., 2023b) and InstructBLIP (Dai et al., 2023) employ Q-Former to connect the frozen LLM and vision encoder. InternVL (Chen et al., 2023c) trains huge ViT and QFormer to integrate visual modalities through a multi-stage training method. MiniGPT4 (Zhu et al., 2023) leverages both a Q-Former and a linear projector to bridge the gap between the vision module and LLM. Furthermore, LLaVA (Liu et al., 2023b) adopts a simple MLP projector to promote the alignment between the LLM and vision encoder. mPLUG-Owl (Ye et al., 2023) introduces an approach that begins to finetune the vision encoder and align visual features, followed by tuning the LLM using LoRA (Hu et al., 2021). Qwen-VL (Bai et al., 2023b) improves visual module resolution to 448, aiming to refine the model’s visual processing capabilities. Fuyu-8B (Bavishi et al., 2023) directly projects image patches before integration with LLM. MM1 (McKinzie et al., 2024) has conducted ablative studies on connector design choices, revealing that the modality adapter type is less critical than the number of visual tokens and the resolution. MiniGemini (Li et al., 2024d) utilizes high-resolution visual tokens and high-quality data to narrow the performance gap with GPT-4 and Gemini. With the rapid advancements in open-source models, proprietary models such as GPT-4o (OpenAI, 2024), Gemini (Team et al., 2023; Reid et al., 2024), Qwen-VL-series (Wang et al., 2024), and Claude3 (anthropic, 2024) have achieved outstanding results in evaluations and practical applications. Some other recent works (Lu et al., 2024b; Zhu et al., 2025; Sun et al., 2025b; Li et al., 2025; Sun et al., 2025a; Dong et al., 2025; Zhang et al., 2024c) provide valuable insights and future directions for building and understanding vision-language models. In this work, owing to the simplicity of the LLaVA architecture, we adopt a framework similar to LLaVA to design our model.
Multilingual Multimodal Models. Recent years have witnessed rapid progress in the expansion of multimodal models to include a wider variety of languages. M3(Ni et al., 2021) leverages English as a pivot and alternates between English-only vision-language pre-training and multilingual masked language modeling. In contrast, UC2 (Zhou et al., 2021) translates English captions into various languages and uses images as the anchor. mCLIP (Chen et al., 2023a) enhances the CLIP model by aligning it with a multilingual text encoder through knowledge distillation. Thanks to the expansion of the overall capabilities of large language models (AI, 2024; Bai et al., 2023a; Jiang et al., 2023; Young et al., 2024), their multilingual capacities have significantly improved. Integrating multilingual LLMs with visual abilities has increasingly become a research focus. In the domain of LLMs, PaLI (Chen et al., 2022) develops a 17B multilingual language-image model that spans over 100 languages. Ying-VLM (Li et al., 2023c) discovers that instruction tuning in English can extend its applicability to other languages. Ziya-Visual (Lu et al., 2023) illustrates the translation of English image-text datasets into Chinese, using in-context learning for instruction-response generation. VisCPM (Hu et al., 2023) introduces a training paradigm that fine-tunes the MLLM in a quasi-zero-shot manner based on a strong multilingual large language model. Despite these advancements, they are primarily confined to two languages or rely on the massive translated corpus. On the other hand, there is no suitable multilingual benchmark for MLLMs to evaluate the performance of multiple languages. There are also some multilingual research studies in other domains, such as multilingual machine translation (Zhao et al., 2024; Pires et al., 2023; Purason & Tättar, 2022; Zhang et al., 2021).
Table 3: Details on the Parrot’s training data, which is derived from publicly available datasets.
Training Stage Datasets Samples Total
Stage 1 LLaVA-1.5-pretrain (Liu et al., 2023b) 558K 1.2M
Laion-Caption (Schuhmann et al., 2022) 12K
CC12M-Caption (Changpinyo et al., 2021) 645K
Stage 2 LLaVA-1.5-finetune (Liu et al., 2023b) 665K 793K
ShareGPT4V-zh (Chen et al., 2023b) 71K
ShareGPT4V-pt (Chen et al., 2023b) 14K
ShareGPT4V-ar (Chen et al., 2023b) 12K
ShareGPT4V-tr (Chen et al., 2023b) 17K
ShareGPT4V-ru (Chen et al., 2023b) 14K
Appendix C Additional Experimental Results
In this section, we present additional experiments and ablation studies to further validate the generality and capability of Parrot across various tasks. Additionally, we elaborate on the training details for Figure 1 to provide a clearer understanding.
C.1 Bilingual Evaluation on LLaVA-Bench
VisCPM (Hu et al., 2023) extends the LLaVA-Bench dataset to the Chinese version for bilingual evaluation. To comprehensively compare Parrot with other multilingual models, we conduct experiments on this benchmark. Due to the deprecation of the GPT-4-0314 version by OpenAI, we test Parrot in LLaVA-Bench following the version of GPT-4-1106-preview for comparison. As shown in Table 4, Parrot not only demonstrates exceptional ability in the English version of this benchmark but also presents competitive performance in the Chinese version.
Notably, as shown in Table 5, VisCPM uses 140M English data and 1M Chinese data to train the model. In comparison, Qwen-VL-Chat uses 1.1B English data and 300M Chinese data, whereas Parrot only utilizes approximately 2M data in total. Despite using less than 1% of the training data, Parrot achieves remarkable performance in both the English and Chinese versions on LLaVA-Bench. Owing to the architecture we proposed, significant improvement in the model’s multilingual capability can be achieved with minimal data usage.
Table 4: Experimental results on LLaVA Test Set accessed by GPT-4. Con: Conversation, DD: Detailed Description, CR: Complex Reasoning, AVG: the average score of three tasks. The symbol denotes that the data are judged following the version of GPT-4-1106-preview because the GPT-4-0314 version is deprecated by OpenAI.
Model LLM Backbone English Chinese
Con DD CR AVG Con DD CR AVG
English Model MiniGPT-4 Vicuna-13B 65.0 67.3 76.6 69.7 - - - -
InstructBLIP Vicuna-13B 81.9 68.0 91.2 80.5 - - - -
LLaVA Vicuna-13B 89.5 70.4 96.2 85.6 - - - -
VisualGLM ChatGLM-6B 62.4 63.0 80.6 68.7 76.6 87.8 83.6 82.7
Qwen-VL-Chat Qwen-7B 82.4 76.9 91.9 83.8 82.3 93.4 89.5 88.2
VisCPM-Balance CPM-Bee-10B 75.5 64.7 91.3 77.3 85.4 81.4 96.6 88.0
Multilingual
Model
Parrot Qwen1.5-7B 82.5 71.0 89.3 81.1 82.1 88.6 92.3 87.7
Table 5: Comparison of vision encoders, LLMs, and training data in different models.
Model vision encoder LLM Training Data
mPLUG-Owl ViT-L/14 (0.3B) BLOOMZ-7B -
VisualGLM Q-Former (1.6B) ChatGLM-6B English: 300M; Chinese 30M
Qwen-VL-Chat ViT-bigG (1.9B) Qwen-7B English: 1.1B; Chinese: 300M
VisCPM Muffin (0.7B) CPM-Bee-10B English: 140M; Chinese: 1M
Parrot ViT-L/14 (0.3B) Qwen1.5-Chat-7B English: 1.8M; Chinese: 71K
C.2 Further Ablation Studies
Ablation study on each component. We conduct an ablation experiment on the multilingual data and the MoE module. As shown in Figure 7, using multilingual data improves performance in each language. Moreover, the MoE module significantly improves performance, demonstrating the effectiveness of our proposed method.
Ablation study on monolingual fine-tuning datasets. The ablation study presented in Table 14 evaluates the performance of different monolingual datasets added incrementally to the baseline dataset LLaVA-1.5-finetune. It highlights the significant impact of adding different multilingual datasets to a baseline model. Each dataset incrementally improves performance in its respective language and, when combined, leads to overall enhanced performance across all evaluated languages. This indicates the robustness and effectiveness of the proposed method in handling multilingual data, making it a scalable solution for multilingual tasks.
Dataset English Chinese Portuguese Arabic Turkish Russian
LLaVA-1.5-finetune 69.4 66.6 60.3 55.3 52.1 60.7
+ zh 69.2 -0.2   68.6 +2.0   64.1 +3.8   59.1 +3.8   50.9 -1.2   61.6 +0.9
+ zh pt 71.1 +1.7   70.4 +3.8   65.4 +5.1   57.9 +2.6   52.1 +0.0 62.9 +2.2
+ zh pt ar 71.0 +1.6   68.6 +2.0   65.7 +5.4   58.6 +3.3   52.2 +0.1   62.2 +1.5
+ zh pt ar tr 70.4 +1.0   68.7 +2.1   64.9 +4.6   61.2 +5.9   59.7 +7.6   62.0 +1.3
+ zh pt ar tr ru 70.7 +1.3   70.4 +3.8   65.1 +4.8   57.8 +2.5   58.4 +6.3   64.0 +3.3
C.3 Comparison of Different Vision Encoders
We also compare the different vision encoders within the Parrot framework in Table 7. It shows that the Chinese-CLIP-based model maintains comparable multilingual performance to the OpenAI-CLIP-based one. This demonstrates that our framework can be compatible with different vision encoders and achieve multilingual alignment through the MoE module.
Table 7: The comparison of various vision encoders within the Parrot framework.
Method LLM Vision Encoder MMMB MMBench
en zh pt ar tr ru en zh pt ar tr ru
LLaVA-1.5 Vicuna-v1.5-7B OpenAI-CLIP 67.07 58.83 59.76 43.50 46.43 59.06 65.37 58.33 59.02 36.16 43.90 56.95
LLaVA-1.5 Vicuna-v1.5-7B Chinese-CLIP 66.45 59.23 59.22 42.68 46.11 58.89 65.92 57.85 58.45 36.90 44.82 56.32
ShareGPT4V Vicuna-v1.5-7B OpenAI-CLIP 69.24 60.23 60.29 43.57 45.26 61.23 69.59 61.60 59.62 37.37 43.38 59.45
ShareGPT4V Vicuna-v1.5-7B Chinese-CLIP 68.65 60.85 59.49 44.33 44.90 61.88 70.28 61.91 58.83 37.00 42.55 58.97
Parrot Qwen1.5-7B OpenAI-CLIP 70.00 68.13 67.31 62.69 58.01 66.26 70.70 70.36 65.12 57.82 58.43 64.00
Parrot Qwen1.5-7B Chinese-CLIP 69.22 69.24 66.32 62.15 57.77 64.31 69.95 70.87 64.92 56.57 57.13 63.15
Method Vision encoder LLM MMB-EN MMB-CN MMMB-en MMMB-zh
LLaVA OpenAI-CLIP ViT-L/14 Vicuna 7B 65.4 58.3 67.1 58.8
LLaVA OpenAI-CLIP ViT-L/14 Qwen1.5-Chat 7B 68.8 66.4 68.2 62.4
LLaVA Chinese-CLIP ViT-L/14 Qwen1.5-Chat 7B 68.1 68.3 67.6 66.1
Parrot OpenAI-CLIP ViT-L/14 Qwen1.5-Chat 7B 70.7 70.4 70.0 68.1
C.4 Comparison with LLaVA using the Same Data
To validate the effectiveness of our proposed approach, we conduct further experiments with an ablation study. Specifically, we expand the baseline LLaVA method by incorporating the same multilingual data used in Parrot. Both models are evaluated on the MMMB dataset, and the results are presented in the Table 10. From the results, we observe that while LLaVA shows a slight improvement with the addition of multilingual data, the increase in performance is limited. In contrast, our Parrot model demonstrates a substantial improvement when multilingual data is included, significantly outperforming LLaVA. This highlights that simply adding multilingual data is not sufficient to bridge the multilingual gap, further emphasizing the effectiveness of our proposed design.
Table 9: The comparison of the baseline LLaVA and Parrot using the same multilingual training data. LLaVA shows limited improvement with multilingual data, while Parrot achieves significant gains, greatly outperforming LLaVA.
Method MMMB
en zh pt ar tr ru
LLaVA w/o Multilingual data 67.1 58.8 59.8 43.5 46.4 59.1
LLaVA w/ Multilingual data 67.0 59.1 60.3 44.2 48.1 59.7
Parrot 70.0 68.1 67.3 62.7 58.0 66.3
Table 10: The comparison of the translation-based baseline and Parrot. While the naive baseline shows some improvements in certain languages, such as Chinese, it leads to performance degradation in others, such as Russian and Portuguese.
Method MMMB
en zh pt ar tr ru
LLaVA 67.1 58.8 59.8 43.5 46.4 59.1
LLaVA w/ translation 67.1 60.7 58.6 47.3 48.6 58.9
Parrot 70.0 68.1 67.3 62.7 58.0 66.3
C.5 Data Scaling and Model Size Scaling
To further investigate the scaling law in multilingual settings, we have conducted experiments where we progressively expanded the multilingual data (excluding Chinese and English) until it reached a volume comparable to the amount of Chinese data ($\sim$70K). The results, shown in the Table 12, demonstrate that Parrot still satisfies the multilingual scaling law. For instance, the performance on Portuguese improved by 3.0 points, and Arabic saw a gain of 5.2 points. As we increase the multilingual data, the model’s performance on the MMMB benchmark continues to improve, suggesting that our model can handle imbalanced multilingual data while still achieving effective scaling and performance gains.
Table 11: The performance comparison on MMMB when scaling the sample sizes of each language.
Sample Size MMMB
en zh pt ar tr ru
10K 70.0 68.1 67.3 62.7 58.0 66.3
30K 70.1 68.0 67.6 64.1 59.9 66.7
50K 69.9 67.9 67.8 64.8 61.4 67.2
70K 70.3 68.4 68.3 65.7 63.2 67.4
Table 12: The performance comparison on MMMB when scaling model sizes of Qwen-series LLM.
Method MMMB
en zh pt ar tr ru
Parrot-7B 70.0 68.1 67.3 62.7 58.0 66.3
Parrot-14B 73.9 71.6 69.8 68.1 64.3 70.1
Parrot-32B 76.3 75.4 73.8 72.1 71.2 73.5
Additionally, we extend Parrot’s LLM backbone from Qwen1.5-7B to Qwen1.5-32B, using the same model design and configuration, and evaluate them on the MMMB dataset. As shown in Table 12, the results indicate that Parrot continues to yield better performance even with a larger LLM backbone. This finding validates the idea that the scaling law for model parameters still holds, and our design remains effective as the model size increases. While we are currently limited to the Qwen1.5-32B model, these results suggest that our approach can scale well with model size, and we believe similar trends would be observed with even larger models, such as those with 30B parameters or beyond.
Appendix D More Implementation Details
D.1 MoE Training Strategy
During the first pre-training stage, the MoE module is initialized with random parameters but is not activated or included in the training process. Instead, we focus exclusively on training the projector. This avoids the issue of training a good projector under a randomly initialized MoE. In detail:
1) Pre-training Stage. In this stage, the MoE module is bypassed entirely, meaning the image tokens do not pass through the MoE. The primary goal of this stage is to train the projector using a large number of image-text pairs. This enables the projector to align image tokens and textual tokens effectively without interference from the untrained MoE module.
2) SFT Stage. Since the SFT stage requires the participation of MoE modules, we randomly initialize the parameters of the MoE components prior to the SFT phase. Once the projector has been trained and achieves robust alignment capabilities in the pre-training stage, we introduce multilingual training data and activate the MoE parameters. At this stage, the MoE is optimized with textual guidance, which drives the alignment of visual tokens while leveraging the well-trained projector. The prior alignment achieved in the pre-training stage allows the MoE to optimize efficiently during this phase.
We present the entire training process of Parrot in the form of pseudocode, as shown in Algorithm 1. It is clear from the algorithm that during the pre-training phase, only the projector is trained. Before the start of the SFT phase, the MoE modules are randomly initialized and incorporated into the training process during the SFT phase.
Algorithm 1 Parrot for multilingual MLLM
Input: Pre-training datasets: $\mathcal{D}^{1}$, SFT datasets: $\mathcal{D}^{2}$;
[…]
D.2 More Experimental Details about Different Backbones
In the following, we provide detailed information to explain Figure 1. Firstly, to ensure a fair comparison between the OpenAI-CLIP-based model and the Chinese-CLIP-based model, we train distinct models using the same training data as LLaVA, as shown in Table 2(a). The hyperparameters are listed in Table 13 without the MoE hyperparameters. As depicted in Figure 1, the OpenAI-CLIP-based model struggles to generate Chinese outputs when given Chinese prompts due to the English-centric training data. In contrast, despite the extremely scarce amount of Chinese training data, the Chinese-CLIP-based model naturally acquires zero-shot capability to understand, process, and generate Chinese texts. Furthermore, we compare both models on MMBench-CN and MMMB-zh to evaluate their Chinese capability. As shown in Table 8, the performance of the Chinese-CLIP-based model is significantly higher than that of the OpenAI-CLIP-based model. On the other hand, we empirically find that different LLMs have a significant impact on performance. Qwen (Bai et al., 2023a) demonstrates superior Chinese capability compared to Vicuna (Chiang et al., 2023), yet its English capability remains competitive.
D.3 Implementation Details of Parrot